Two sides of an equilateral triangle have lengths 2x - 2 and 3x - 6. Which could be the length of the third side: 10 - x or 6x - 5?
Answer by richard1234(7193)

It is necessary that 2x-2 and 3x-6 are equal since the triangle is equilateral, so 2x-2 = 3x-6 --> x = 4. The side length of the triangle is 2*4 - 2 = 6. Plugging in x = 4, 10-x = 6 and 6x-5 = 17, so 10-x must be the third side.